# Measurement of the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

### v1

@@ III.2.1 Four-particle symmetric cumulants
Next, we compare the results from our simulations with the experimental measurements from RHIC. $SC(m,n)$ have been measured in previous studies from hydrodynamic and transport models [44, 43, 37, 45, 46, 47, 48], while $AC_{2,1}(m,n)$ have been measured in some transport models [46, 47]. Fig. [3] shows comparisons between the theory calculations for the symmetric cumulants $SC(2,3)$ and $SC(2,4)$ multiplied by $\langle N_{part}\rangle$ and the same measured by the STAR Collaboration [71]. The factor of $\langle N_{part}\rangle$ is multiplied to scale out the trivial dilution of correlation with the increase in the number of triplets. We are able to reproduce the qualitative variations with centrality. Here, we don’t aim to reproduce the experimental data because our aim is to check the sensitivity of these observables to the hydro transport coefficients. The comparison in Fig.[3] reflects that the correlation in initial spatial eccentricities and the following hydrodynamic evolution can capture the correlation among flow harmonics of different orders. Negative values of $SC(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as $SC(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. Symmetric cumulants do not have contributions from non-flow effects, where non-flow refers to azimuthal correlations not related to the reaction plane orientation, like those from resonances, jets, quantum statistics, etc. This is verified by computing these observables for HIJING model, which includes only non-flow physics, for which these are consistently zero [35]. The model captures the variation of $SC(m,n)$ with the centrality. In central collisions, the model underestimates $SC(2,4)$, possibly due to inadequate fluctuations in the initial conditions or some shortcomings in the transport model since the nonlinear response in $v_{4}$ and $\varepsilon_{4}$ is sensitive to these effects. However, $SC(2,3)$ is better aligned with the data because the $v_{2}-v_{3}$ correlations mainly arise from the initial state $\varepsilon_{2}-\varepsilon_{3}$ correlations due to linear response in $v_{2}$ and $v_{3}$.
@@
Figure 3: Symmetric cumulant vs centrality from four-particle correlations compared with STAR data [71] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. $SC(2,3)$ is consistently negative while as $SC(2,4)$ is consistently positive.
The normalized symmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV, evaluated using the Monte Carlo Glauber model in coordinate space (using Eq.[11]), are presented in Fig.[4] and compared with $NSC(m,n)$ measured in momentum space. If only eccentricity drives $v_{n}$, then we can expect that the $NSC(m,n)$ in the final state would be equal to the $NSC(m,n)$ initial state. Fig.[4] shows that the initial anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ is mainly responsible for the observed anti-correlation between $v_{2}$ and $v_{3}$. However, the correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is smaller than the observed correlation between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ not only comes from the linear response of the system to $\varepsilon_{4}$ but also has a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between $NSC(2,4)$ in the final state and the initial state increases, likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$. This has also been observed in Pb+Pb collisions at $\sqrt{s_{NN}}$ = 2.76 TeV by the ATLAS [72] and ALICE [39] experiments. The properties of the medium were suggested to affect the relative contribution of $\varepsilon_{2}$ in $v_{4}$ compared to $\varepsilon_{4}$ [73]. Therefore, $NSC(2,4)$ provides a probe into the medium properties.
@@
Figure 4: Normalized symmetric cumulant vs centrality from four-particle correlations compared with STAR data [71] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. $NSC(2,3)$ is consistently negative while as $NSC(2,4)$ is consistently positive.
III.2.2 Six-particle asymmetric cumulants
$AC_{2,1}(m,n)$ have measured by the ALICE collaboration for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [49]. In Fig.[5], we should the centrality dependence of $AC_{2,1}(m,n)$ for Au+Au at $\sqrt{s_{NN}}=200\ GeV$. The trends are similar to the results by the ALICE collaboration. As for symmetric cumulants, negative values of $AC(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as $AC(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. To check their sensitivity to the hydro transport coefficients, we compute these observables for the three sets of ensembles mentioned in Table  2. We observe that both $AC_{2,1}\{2,3\}$ and $AC_{2,1}\{2,4\}$ are suppressed by shear and bulk viscosities. For a clearer picture, the same results are plotted with a log scale for the y-axis in Fig.[6].
Figure 5: Asymmetric cumulant, $AC_{2,1}\{2,3\}$/$AC_{2,1}\{2,4\}$, vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV.
Figure 6: Asymmetric cumulants, -$AC_{2,1}\{2,3\}$ and $AC_{2,1}\{2,4\}$, vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The upper panel shows $AC_{2,1}\{2,3\}$ multliplied by -1 to show the y-axis in log scale while as the lower panel shows $AC_{2,1}\{2,4\}$.
III.3 Conclusions
@@
First, we adjusted the free parameters, such as shear and bulk viscosities, to accurately describe particle multiplicities, mean transverse momentum, and anisotropic flow. After that, we have discussed the impact of shear and bulk viscosities on traditional and novel observables.
We have studied correlators that measure the correlations between flow harmonics of varying orders, specifically four and six-particle correlations, including both symmetric and asymmetric cumulants as functions of centrality at midrapidity in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. These new observables manifestly satisfy all mathematical properties of cumulants in the basis in which they are expressed. This new approach allows for the separation of non-flow and flow contributions and provides initial insights into how the combinatorial background contributes to flow measurements using correlation techniques.
The hybrid framework that we have used yields good agreement with the majority of the latest multi-particle observables from RHIC. In addition, the hadronic afterburner plays an essential role in describing the data collected at RHIC [44]. Study based on HIJING demonstrated that these new observables are robust against systematic biases due to non-flow correlations [43].
Compared to traditional flow observables, the novel observables are more sensitive to transport coefficients, resulting in better constraints on the transport properties of the medium created in heavy-ion collisions. We observed anti-correlation between event-by-event fluctuations of $v_{2}$ and $v_{3}$, while the event-by-event fluctuations of $v_{2}$ and $v_{4}$ are found to be positively correlated. The observed anti-correlation between $v_{2}$ and $v_{3}$ appears to be described by the initial-stage anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$, which supports the idea of linearity between $\varepsilon_{n}$ and $v_{n}$ [23]. The hydrodynamic response suggests that the final state fluctuation originates from the initial state. However, including the nonlinear hydrodynamic response of the medium is necessary to reproduce the measured correlation between $v_{2}$ and $v_{4}$, as the initial-stage correlation alone is insufficient. Therefore, we can gain new knowledge regarding the initial conditions and properties of QGP in high-energy nuclear collisions from these higher-order observables. Symmetric and asymmetric cumulants of flow amplitudes can be used directly to constrain the multivariate probability density function of flow fluctuations since its functional form can be reconstructed only from its true moments or cumulants.
III.4 Acknowledgements

### v2

@@ III.2.1 Four-particle symmetric cumulants
The hadronic afterburner plays an essential role in describing the data collected at RHIC [43]. To check their sensitivity to hadronic interactions and resonance decay, we compute these observables at the end of the three stages of the evolution mentioned in Table  3. We observe that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by resonance decay and enhanced by hadronic interactions. However, the magnitudes of these effects are different for the two observables. In Fig.[3], we show the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. In $v_{2}\{4\}$, we see a change of around 15% due to resonance decay and 5% due to hadronic interactions plus resonance decay in the mid-central collisions. While in ${\rm SC}(2,3)$, we see a change of around 40% due to resonance decay throughout the centrality range and 40-60% due to hadronic interactions plus resonance decay. ${\rm SC}(2,4)$ is also affected in the same manner. Thus, these observables are more sensitive to resonance decay and hadronic interactions. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also reveals similar differences in their sensitivity to these effects.
Figure 3: Ratios of $v_{2}\{4\}$ and ${\rm SC}(2,3)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm SC}(2,3)$ using lines and markers. The red line and the cyan band show the effects of resonance decay, and the blue line and the magenta band show the effects of hadronic interactions.
III.2.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig.[4], we show the effects of resonance decay and hadronic interactions on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,4)$. For ${\rm AC}_{2,1}(2,4)$, we see a change of around 40% due to these in the mid-central collisions. Although the mean values are somewhat similar throughout centrality, due to large uncertainties (about 2X), AC results for 0-10% and 10-20% centralities have been excluded for clarity. ${\rm AC}_{2,1}(2,3)$ is also affected in the same manner. Compared to the traditional flow observables, we see that these are more sensitive to resonance decay and hadronic interactions.
Figure 4: Ratios of $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm AC}_{2,1}(2,4)$ using lines and markers. The red line and the cyan band show the effects of resonance decay, and the blue line and the magenta band show the effects of hadronic interactions. AC results for 0-10% and 10-20% centralities, with large uncertainties, have been excluded for clarity.
III.3 Centrality dependence of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$
III.3.1 Four-particle symmetric cumulants
@@
Next, we compare the results from our simulations with the experimental measurements from RHIC. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [43, 42, 37, 44, 45, 46, 47], while ${\rm AC}_{2,1}(m,n)$ have been computed in some transport models [45, 46]. Fig. [5] shows comparisons between the theory calculations for the symmetric cumulants ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ multiplied by $\langle N_{part}\rangle$ and the same measured by the STAR Collaboration [70]. The factor of $\langle N_{part}\rangle$ is multiplied to scale out the trivial dilution of correlation with the increase in the number of multiplets. We are able to reproduce the qualitative variations with centrality. The magnitudes of $\langle N_{part}\rangle{\rm SC}(2,3)$ and $\langle N_{part}\rangle{\rm SC}(2,4)$ increase from central to mid-peripheral events and then again decreases for very peripheral events. Here, we do not aim to reproduce the experimental data because our aim is to check the sensitivity of these observables to the transport properties of the system created in these collisions. The comparison in Fig.[5] reflects that the correlation in initial spatial eccentricities and the following hydrodynamic evolution can capture the correlation among flow harmonics of different orders. Negative values of ${\rm SC}(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as ${\rm SC}(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. Symmetric cumulants do not have contributions from non-flow effects, where non-flow refers to azimuthal correlations not related to the reaction plane orientation, like those from resonances, jets, quantum statistics, etc. This is verified by computing these observables for the HIJING model, which includes only non-flow physics, for which these are consistently zero [35]. The model captures the variation of ${\rm SC}(m,n)$ with the centrality. In central collisions, the model underestimates ${\rm SC}(2,4)$, possibly due to inadequate fluctuations in the initial conditions or some shortcomings in the transport model since the nonlinear response in $v_{4}$ and $\varepsilon_{4}$ is sensitive to these effects. However, ${\rm SC}(2,3)$ is better aligned with the data because the $v_{2}-v_{3}$ correlations mainly arise from the initial state $\varepsilon_{2}-\varepsilon_{3}$ correlations due to linear response in $v_{2}$ and $v_{3}$.
@@
Figure 5: Symmetric cumulant vs centrality from four-particle correlations compared with STAR data [70] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm SC}(2,3)$ is consistently negative while as ${\rm SC}(2,4)$ is consistently positive.
The normalized symmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV, computed using Eq.[5] in momentum space and the corresponding equation in coordinate space for the initial state eccentricity (Eq.[11]) evaluated using the Monte Carlo Glauber model, are presented in Fig.[6].
@@
If only eccentricity drives $v_{n}$, then we can expect that the ${\rm NSC}(m,n)$ in the final state would be equal to the ${\rm NSC}(m,n)$ in the initial state. Fig.[6] shows that the initial anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ is mainly responsible for the observed anti-correlation between $v_{2}$ and $v_{3}$. However, the correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is smaller than the observed correlation between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ not only comes from the linear response of the system to $\varepsilon_{4}$ but also has a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between ${\rm NSC}(2,4)$ in the final state and the initial state increases, likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$. This has also been observed in Pb+Pb collisions at $\sqrt{s_{NN}}$ = 2.76 TeV by the ATLAS [71] and ALICE [39] experiments. The properties of the medium were suggested to affect the relative contribution of $\varepsilon_{2}$ in $v_{4}$ compared to $\varepsilon_{4}$ [72]. Therefore, ${\rm NSC}(2,4)$ provides a probe into the medium properties.
Figure 6: Normalized symmetric cumulant vs centrality from four-particle correlations compared with STAR data [70] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm NSC}(2,3)$ is consistently negative while as ${\rm NSC}(2,4)$ is consistently positive.
III.3.2 Six-particle asymmetric cumulants
${\rm AC}_{2,1}(m,n)$ have measured by the ALICE collaboration for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [48]. In Fig.[7], we should the centrality dependence of ${\rm AC}_{2,1}(m,n)$, multiplied by $\langle N_{part}\rangle$ to scale out the trivial dilution of correlation with the increase in the number of multiplets, for Au+Au at $\sqrt{s_{NN}}=200\ GeV$. The magnitudes of $\langle N_{part}\rangle{\rm AC}_{2,1}(2,3)$ and $\langle N_{part}\rangle{\rm AC}_{2,1}(2,4)$ increase with centrality. The trends are similar to the results by the ALICE collaboration. As for symmetric cumulants, negative values of $AC(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as $AC(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. We compute these observables for the three sets of ensembles mentioned in Table  2 and observe that both ${\rm AC}_{2,1}(2,3)$ and ${\rm AC}_{2,1}(2,4)$ are suppressed by shear and bulk viscosities.
Figure 7: Asymmetric cumulant vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm AC}_{2,1}(2,3)$ is consistently negative while as ${\rm AC}_{2,1}(2,4)$ is consistently positive.
In Fig.[8], we have compared the normalized asymmetric cumulant in coordinate and momentum space, as computed using Eq.[8] and the corresponding equation for the initial state eccentricities (Eq.[11]). Again, if only eccentricity drives $v_{n}$, we can expect the ${\rm NAC}(m,n)$ in the final state to be equal to the ${\rm NAC}(m,n)$ in the initial state. Fig. [8] illustrates that the initial negative correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ primarily causes the observed negative correlation between $v_{2}$ and $v_{3}$. However, unlike symmetric cumulants, here, the initial state correlations are smaller in the peripheral collisions. The correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is weaker than the correlation observed between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ comes not only from the linear response of the system to $\varepsilon_{4}$, but also from a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between ${\rm NAC}(2,4)$ in the final state and the initial state increases. This is likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$.
Figure 8: Normalized asymmetric cumulants vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm NAC}_{2,1}(2,3)$ is consistently negative while as ${\rm NAC}_{2,1}(2,4)$ is consistently positive.
III.4 Conclusions
@@
First, we adjusted the free parameters, such as shear and bulk viscosities, to describe particle multiplicities, mean transverse momentum, and anisotropic flow. After that, we discussed the impact of shear and bulk viscosities, resonance decay, and hadronic interactions on traditional and novel observables.
We have studied correlators that measure the correlations between flow harmonics of varying orders, specifically four and six-particle correlations, including both symmetric and asymmetric cumulants as functions of centrality at midrapidity in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. These observables are very sensitive to shear and bulk viscosities, resonance decay, and hadronic interactions. A single parameter set of transport coefficients is unable to explain $SC(2,3)$ and $SC(2,4)$ simultaneously.
We observed anti-correlation between event-by-event fluctuations of $v_{2}$ and $v_{3}$, while the event-by-event fluctuations of $v_{2}$ and $v_{4}$ are found to be positively correlated. The observed anti-correlation between $v_{2}$ and $v_{3}$ appears to be described by the initial-stage anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$, which supports the idea of linearity between $\varepsilon_{n}$ and $v_{n}$ [23]. The hydrodynamic response suggests that the final state fluctuation originates from the initial state. However, including the nonlinear hydrodynamic response of the medium is necessary to reproduce the measured correlation between $v_{2}$ and $v_{4}$, as the initial-stage linear correlation alone is insufficient.
III.5 Acknowledgements